# Artificial Precision Polarization Array: Sensitivity for the axion-like dark matter with clock satellites

### v2

@@ III.2 Frequentist analysis-GLSP
This section is based on the following assumptions: the cross-correlation between transmitting satellites can be safely neglected. For further details of the analysis, please refer to Appendix B. We employ simulated observational data that incorporates scalar axion field signals and apply the GLSP method to perform frequentist analysis in the search for effective signals.
@@
$P_{LS}\big(f\big)$ $=\frac{1}{\sigma^{2}_{y}\sum_{i}\omega_{i}}\Bigg\{\frac{\big[\sum_{i}\omega_{i}\big(y_{i}-\bar{y}\big)\cos\big(2\pi f\big(t_{i}-\tau\big)\big)\big]^{2}}{\sum_{i}\omega_{i}\cos^{2}\big[2\pi f\big(t_{i}-\tau\big)\big]}$ (III.6)
@@
$\tau=\frac{1}{4\pi f}\arctan\frac{\sum_{i}\omega_{i}\sin\big(4\pi ft_{i}\big)}{\sum_{i}\omega_{i}\cos\big(4\pi ft_{i}\big)},$ (III.7)
@@
$\omega_{i}$ $=\frac{1}{\sigma^{2}_{i}},\quad\bar{y}=\frac{\sum_{i}\omega_{i}y_{i}}{\sum_{i}\omega_{i}},$ (III.8)
@@
$\Delta\phi_{\text{sim}}=\phi_{a}\mathrm{cos}\big(2\pi\nu+\varphi\big)+\phi_{\text{noise}}.$ (III.9)
@@
If the observational data are of low quality and the target frequency peak is completely masked by interference peaks, the GLSP method becomes ineffective. In such cases, for further information, the reader is referred to [43, 12]. Once the GLSP method successfully extracts the target axion signal frequency $m_{a}$ from the pulsed signals received by the satellite, the APPA sensitivity curve for axion signal detection can be derived using Eq. (B.1) and Eq. (III.9).
We assume that the satellite network possesses a detection sensitivity of $0.01^{\circ}$ for linear polarization angle rotation, and that the instrumental noise is parameterized according to Eq. (III.1) and Eq. (III.2). Under these assumptions, the sensitivity curve can be obtained subject to the constraint of a total observation time of $T_{\text{obs}}\sim\mathcal{O}\big(10^{1}\big)$yr. Note that when the axion mass is sufficiently small ($\mathrm{cos}\big(m_{a}T\big)\simeq 1$, to the left of the dashed line in Fig. 2), Eq. (B.1) reduces to Eq. (B.2), in this limit, the expected value of $\left|\alpha_{o}-\alpha_{s}\right|$ is 0.734; when condition $\mathrm{cos}\big(m_{a}T\big)\simeq 1$ is no longer satisfied (to the right of the dashed line in Fig. 2, we adopt $m_{a}L=0.1$ as the reference criterion), the expectation value of $\big[\big(\alpha_{o}^{2}+\alpha_{s}^{2}\big)-2\alpha_{o}\alpha_{s}\mathrm{cos}\big(m_{a}T\big)\big]^{\frac{1}{2}}$ must be evaluated individually for each value of $m_{a}$. Unlike the likelihood analysis, the spatial scale of the satellite network only determines the critical line at which $\mathrm{cos}\big(m_{a}T\big)\simeq 1$ and does not affect the overall sensitivity curve. Detailed analysis can be found in Appendix B.
@@ IV Conclusion
To clarify the feasibility and advantages of APPA, we employ two complementary approaches: likelihood analysis and frequentist analysis. For the likelihood analysis, we avoid prohibitively large scale computations by using the convenient expression in Eq. (III.4) from [14, 19], which allows us to efficiently derive the $g_{95\%}$ constraint. Likelihood analysis indicates that larger scales $L$ of satellite networks are more advantageous for detecting low mass axions. However, for axions with $\lambda_{\text{com}}\geq L$, APPA exhibit identical detection limits. Fig. 2 shows that APPA exhibits a detection advantage within $m_{a}\sim\mathcal{O}\big(10^{-22}–10^{-18}\big)$eV, which lies well below the detection upper bound. Compared with traditional ground-based observations, APPA achieves a tighter $g_{95\%}$ bound, highlighting its strong detection potential. For the frequentist approach, we adopt the GLSP, a standard method for detecting periodic sinusoidal signals. MC simulations demonstrate that, even under low S/N conditions, APPA produces GLSP spectra with more pronounced target signal peaks, significantly reduced FAP, and fewer interference peaks compared with conventional observations.
@@ Appendix B GLSP: Generalized Lomb-Scargle Periodogram
Under the APPA framework, the satellite network is confined to a small scale region of the Solar System with $L_{i}\approx 5$AU (Jupiter’s orbit). For axions velocity $v\sim 10^{-3}c$, the de Broglie wavelength $\lambda_{\text{de}}$ of the scalar axion field is $\lambda_{\text{de}}=\frac{h}{m_{a}v}\sim\mathcal{O}\big(10^{-5}-10^{2}\big)\mathrm{pc}$, which is much larger than the spatial extent of the satellite array for axion masses in the range $m_{a}\sim\mathcal{O}\big(10^{-22}-10^{-15}\big)$eV. Therefore, the entire network can be treated as residing within a single coherent axion field, implying $\varphi_{o}=\varphi_{s}=\varphi$ and $\rho_{o}=\rho_{s}=\rho_{\text{DM}}$. Consequently, a significant fraction of environmental disturbances and intrinsic signal uncertainties are effectively eliminated, and Eq. (II.10) reduces to
$\phi_{a}$ $=\frac{g_{a\gamma}}{m_{a}}\sqrt{\frac{\rho_{\text{DM}}}{2}}\big[\left(\alpha_{o}^{2}+\alpha_{s}^{2}\right)-2\alpha_{o}\alpha_{s}\mathrm{cos}\left(m_{a}T\right)\big]^{\frac{1}{2}}.$ (B.1)
Figure 6: The histogram (yellow, $bin=100$) of $\left|\alpha_{o}-\alpha_{s}\right|$ and its PDF (blue). Using the MC method, $10^{6}$ samples were generated, yielding the mean value $\mu=0.734$ and the variance $\sigma=0.319$.
Since the satellite network is distributed on a spatial scale of 55AU, the corresponding signal transmission time is $T\sim\mathcal{O}\big(10^{3}\big)$s. Thus, for the axion masses $m_{a}\leq 2.6\times 10^{-20}$eV, one has $\mathrm{cos}\big(m_{a}T\big)\simeq 1$. Under this condition, Eq. (B.1) can be rewritten as
$\phi_{a}$ $\simeq\frac{g_{a\gamma}}{m_{a}}\sqrt{\frac{\rho_{\text{DM}}}{2}}\left|\alpha_{o}-\alpha_{s}\right|.$ (B.2)
Using MC method, we derive the statistical histogram and probability density function (PDF) of $\left|\alpha_{o}-\alpha_{s}\right|$, the results are presented in Fig. 6.
@@
Figure 7: The Histograms and corresponding PDFs of $\big[\left(\alpha_{o}^{2}+\alpha_{s}^{2}\right)-2\alpha_{o}\alpha_{s}\mathrm{cos}\left(m_{a}T\right)\big]^{\frac{1}{2}}$ for different axion masses $m_{a}$. The distributions were generated using the MC method with $10^{6}$ samples for each $m_{a}$. Color coding: Red, $\mathrm{cos}\big(m_{a}T\big)=1$; Orange, $m_{a}=10^{-19}\mathrm{eV}$; Yellow, $m_{a}=10^{-18}\mathrm{eV}$; Green, $m_{a}=10^{-17}\mathrm{eV}$; Blue, $m_{a}=10^{-16}\mathrm{eV}$; Purple, $\mathrm{cos}\big(m_{a}T\big)=-1$. The black curve represents the envelope defined by the maxima of the PDFs across different values of $m_{a}$.
For the axion mass $m_{a}\geq 2.6\times 10^{-20}$eV, Eq. (B.2) no longer applies, the full expression $\big[\left(\alpha_{o}^{2}+\alpha_{s}^{2}\right)-2\alpha_{o}\alpha_{s}\mathrm{cos}\left(m_{a}T\right)\big]^{\frac{1}{2}}$ must be considered. The corresponding distribution results are shown in Fig. 7. Note that for arbitrary axion mass, the statistical distribution of amplitude fluctuations lies between the two limiting cases of $\mathrm{cos}\big(m_{a}T\big)=1$ (red reigon in Fig. 7) and $\mathrm{cos}\big(m_{a}T\big)=-1$ (purple reigon in Fig. 7).
@@
To further simplify the analysis, we demonstrate that, when processing APPA data with frequentist analysis, the cross-correlation between transmitting satellites can be safely neglected. As an illustrative example, taking $\mathrm{cos}(m_{a}T)=1$, we define the correlation coefficients as follows:
@@
$P_{ab}$ $=\overline{\Delta\phi_{a}(t)\Delta\phi_{b}(t)}=\left(\frac{g_{a\gamma}}{m_{a}}\sqrt{\frac{\rho_{\text{DM}}}{2}}\right)^{2}\Big\langle\left|\alpha_{o}-\alpha_{sa}\right|$
$\quad\times\cos\left(m_{a}t+\varphi_{a}\right)\left|\alpha_{o}-\alpha_{sb}\right|\cos\left(m_{a}t+\varphi_{b}\right)\Big\rangle$
$\propto\frac{1}{2}\big\langle\left|\alpha_{o}-\alpha_{sa}\right|\left|\alpha_{o}-\alpha_{sb}\right|\cos\big(\varphi_{a}-\varphi_{b}\big)\big\rangle.$ (B.3)
@@
In the derivation, since the observation time $T_{\text{obs}}\sim\mathcal{O}\big(10^{1}\big)\mathrm{yr}$ is extremely long, the time averaged value of the high frequency oscillatory term $\cos\big(2m_{a}t+\varphi_{a}+\varphi_{b}\big)$ over the observation period is effectively zero. To estimate the correlation coefficients $P_{ab}$, a total of $10^{6}$ samples were generated using the MC method, and the PDF of $P_{ab}$ is shown in Fig. 8. The results indicate that the correlation coefficients between different transmitting satellites are strongly clustered around zero. Therefore, when analyzing APPA data, the inter signal correlations between distinct satellites can be safely neglected.
@@
Figure 8: The histogram of Eq. (B) (yellow, 100 bins) together with its PDF (blue) is shown. A total of $10^{6}$ samples were generated using the MC method.

### v3

@@ III.2 Frequentist analysis-GLSP
In this section, we employ simulated observational data that incorporates scalar axion field signals and analyse the detection sensitivity of APPA based on the Generalized Lomb–Scargle Periodogram (GLSP) method. Our research is based on the following assumptions: the cross-correlation between transmitting satellites can be safely neglected. For further details of the analysis, please refer to Appendix B.
@@
$P_{LS}\big(f\big)$ $=\frac{1}{\sigma^{2}_{y}\sum_{i}\omega_{i}}\Bigg\{\frac{\big[\sum_{i}\omega_{i}\big(y_{i}-\bar{y}\big)\cos\big(2\pi f\big(t_{i}-\tau\big)\big)\big]^{2}}{\sum_{i}\omega_{i}\cos^{2}\big[2\pi f\big(t_{i}-\tau\big)\big]}$ (III.10)
@@
$\tau=\frac{1}{4\pi f}\arctan\frac{\sum_{i}\omega_{i}\sin\big(4\pi ft_{i}\big)}{\sum_{i}\omega_{i}\cos\big(4\pi ft_{i}\big)},$ (III.11)
@@
$\omega_{i}$ $=\frac{1}{\sigma^{2}_{i}},\quad\bar{y}=\frac{\sum_{i}\omega_{i}y_{i}}{\sum_{i}\omega_{i}},$ (III.12)
@@
$\Delta\phi_{\text{sim}}=\phi_{a}\mathrm{cos}\big(2\pi\nu t+\varphi\big)+\phi_{\text{noise}}.$ (III.13)
@@
If the observational data are of low quality and the target frequency peak is completely masked by interference peaks, the GLSP method becomes ineffective. In such cases, for further information, the reader is referred to [46, 14]. Once the GLSP method successfully extracts the target axion signal frequency $m_{a}$ from the pulsed signals received by the satellite, the APPA sensitivity curve for axion signal detection can be derived using Eq. (III.1) and Eq. (III.13).
We assume that the satellite network possesses a detection sensitivity of $0.003^{\circ}$ for linear polarization angle rotation, and that the instrumental noise is parameterized according to Eq. (III.2) and Eq. (III.3). Under these assumptions, the sensitivity curve can be obtained subject to the constraint of a total observation time of $T_{\text{obs}}\sim\mathcal{O}\big(10^{1}\big)$ yr with Eq. (III.1) and Eq. (III.13), the analysis result is shown in Fig. 6. Detailed analysis can be found in Appendix B.
Figure 6: Sensitivity curve of the APPA satellite network with frequentist analysis. The detection sensitivity is set to $\phi_{a}=0.003^{\circ}$, with minor fluctuations attributed to instrumental noise. The black dashed line marks the critical boundary, to the left of this boundary, the condition $\mathrm{cos}\big(m_{a}T\big)\simeq 1$ holds true, while to the right it is no longer valid (we adopt $m_{a}T=0.1$ as the reference criterion).
@@ IV Conclusion
To clarify the feasibility and advantages of APPA, we employ two complementary approaches: likelihood analysis and frequentist analysis. For the likelihood analysis, we avoid prohibitively large scale computations by using the convenient expression in Eq. (III.6) from [16, 21], which allows us to efficiently derive the $g_{95\%}$ constraint. Likelihood analysis indicates that larger scales $L$ of satellite networks are more advantageous for detecting low mass axions. However, APPA exhibits identical detection limits when axions satisfy $\lambda_{\text{c}}\geq L$, for different $L$. Fig. 2 shows that APPA exhibits a detection advantage within $m_{a}\sim\mathcal{O}\big(10^{-22}–10^{-18}\big)$ eV, which lies well below the detection upper bound. Compared with traditional ground-based observations, APPA achieves a tighter $g_{95\%}$ bound, highlighting its strong detection potential. For the frequentist approach, we adopt the GLSP, a standard method for detecting periodic sinusoidal signals. MC simulations demonstrate that, even under low S/N conditions, APPA produces GLSP spectra with more pronounced target signal peaks, significantly reduced FAP, and fewer interference peaks compared with conventional observations.
@@ Appendix B GLSP: Generalized Lomb-Scargle Periodogram
Under the APPA framework, the satellite network is confined to a small scale region of the Solar System with $L_{i}\approx 5$AU (Jupiter’s orbit). As mentioned in the main text, for axions velocity $v\sim 10^{-3}c$, the de Broglie wavelength $\lambda_{\text{de}}$ of the scalar axion field is larger than the scale of satellite networks. Therefore, the entire network can be treated as residing within a single coherent axion field, then Eq. (II.11) reduces to Eq. (III.1).
@@
Figure 7: The Histograms and corresponding PDFs of $\alpha\big[2\left(1-\mathrm{cos}\left(m_{a}T\right)\right)\big]^{\frac{1}{2}}$ for different axion masses $m_{a}$, under the condition $\alpha_{o}=\alpha_{s}=\alpha$. The distributions were generated using the MC method with $10^{6}$ samples for each $m_{a}$. Color coding: Red, $m_{a}=10^{-19}\mathrm{eV}$; Green, $m_{a}=10^{-17.5}\mathrm{eV}$; Blue, $\mathrm{cos}\big(m_{a}T\big)=-1$. The black curve represents the envelope defined by the maxima of the PDFs across different values of $m_{a}$, which diverges at the end when approaching zero that $\mathrm{cos}\big(m_{a}T\big)=1$.
Using MC method, we derive the statistical histogram and probability density function (PDF) of Eq. (III.1) for different $m_{a}$, the results are presented in Fig. 7. Note that for arbitrary $m_{a}$, the statistical distribution of amplitude fluctuations lies between the two limiting cases of $\mathrm{cos}\big(m_{a}T\big)=1$ and $\mathrm{cos}\big(m_{a}T\big)=-1$ (see Fig. 7).
@@
To further simplify the analysis, we demonstrate that when processing APPA data with frequentist analysis, the cross-correlation between transmitting satellites can be safely neglected. Taking practical orbital factors into account, the distances between the transmitting satellites (a,b=1,,6a,b=1,...,6) and the receiving satellite are changing with time during the operation of the satellite network (as in Fig. 1). According to Eq. (III.1), the signal transmission time $T_{a}$ and $T_{b}$ vary continuously, which results in the randomness of $\varphi_{a}$ and $\varphi_{b}$. Using Eqs. (II.10) and (III.1), we obtain the correlation coefficient as follows:
$P_{ab}$ $=\overline{\Delta\phi_{a}(t)\Delta\phi_{b}(t)}$ (B.1)
@@
$=2\left(\frac{g_{a\gamma}}{m_{a}}\sqrt{\frac{\rho_{\text{DM}}}{2}}\right)^{2}\alpha^{2}\Big\langle\big[\big(1-\cos m_{a}T_{a}\big)\big(1-\cos m_{a}T_{b}\big)\big]^{\frac{1}{2}}$
$\quad\times\cos\left(m_{a}t+\varphi_{a}\right)\cos\left(m_{a}t+\varphi_{b}\right)\Big\rangle$
$\propto\alpha^{2}\Big\langle\big[\big(1-\cos(m_{a}T_{a})\big)\big(1-\cos(m_{a}T_{b})\big)\big]^{\frac{1}{2}}$
$\quad\times\cos\big(\varphi_{a}-\varphi_{b}\big)\Big\rangle.$
@@
In the derivation, since the observation time $T_{\text{obs}}\sim\mathcal{O}\big(10^{1}\big)\mathrm{yr}$ is extremely long, the time averaged value of the high frequency oscillatory term $\cos\big(2m_{a}t+\varphi_{a}+\varphi_{b}\big)$ over the observation period is effectively zero. To estimate the correlation coefficients $P_{ab}$, a total of $10^{6}$ samples were generated using the MC method, and the PDF of $P_{ab}$ is shown in Fig. 8. The results indicate that the correlation coefficients between different transmitting satellites are clustered around zero. Therefore, when analyzing APPA data with GLSP method, the inter signal correlations between distinct satellites can be neglected (see e.g. [46]).
@@
Figure 8: The histogram of Eq. (B.1) (yellow, 100 bins) together with its PDF (blue) is shown. A total of $10^{6}$ samples were generated using the MC method.